# MusRec: Zero-Shot Text-to-Music Editing via Rectified Flow and Diffusion Transformers

### v1

@@ Abstract
Music editing has emerged as an important and practical area of artificial intelligence, with applications ranging from video game and film music production to personalizing existing tracks according to user preferences. However, existing models face significant limitations, such as being restricted to editing synthesized music generated by their own models, requiring highly precise prompts, or necessitating task-specific retraining—thus lacking true zero-shot capability.
Leveraging recent advances in rectified flow and diffusion transformers, we introduce MusRec, the first zero-shot text-to-music editing model capable of performing diverse editing tasks on real-world music efficiently and effectively. Experimental results demonstrate that our approach outperforms existing methods in preserving musical content, structural consistency, and editing fidelity, establishing a strong foundation for controllable music editing in real-world scenarios.
@@ I Introduction
At the same time, a parallel line of research has introduced rectified flow models [lee2024improvingtrainingrectifiedflows, liu2022flowstraightfastlearning], which reformulate diffusion as a more direct flow between noise and data distributions. Rectified flows enable efficient and stable generation, and have recently been realized at scale through the Flux family of diffusion transformers [tang2024hart, xie2024sana, yang2024cogvideox, peebles2023scalable]. Flux Plays Music in particular demonstrated the power of this approach for text-to-music generation. In computer vision, the work of Taming Rectified Flow for Inversion and Editing [fluxplaysmusic] showed that RF models also support accurate inversion and robust editing, but these ideas have not yet been applied to music. This raises an intriguing question: can the strengths of rectified flow be used not just for generating music, but for editing real recordings in a practical, zero-shot fashion?
@@ I-B Contributions
  • We present the first framework for rectified-flow based editing of real music recordings.
@@ II-B Music Editing
In this work, we introduce MusRec, a zero-shot framework built on pretrained rectified flow models for music editing. MusRec injects the self-attention features of the source music from a diffusion transformer directly into the editing process. Unlike prior approaches, it can operate effectively on real-world audio and generalizes across a wide range of editing tasks. Moreover, MusRec removes the need for prompt engineering during both reversal and editing, making music editing more accessible and practical.
@@ IV-B RF-Solver
$+\tfrac{1}{2}(t_{i-1}-t_{i})^{2}\,\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (8)
@@ IV-C Attention Feature Replacement Strategies
To explore the effect of feature-level guidance during the denoising process, we further experimented with several attention feature replacement strategies within the diffusion transformer architecture. The goal of these experiments is to investigate how reusing intermediate representations from the inversion stage can improve controllability and structure preservation in the generated outputs. Inspired by prior work on attention-level feature reuse in rectified flow models [tamingrectifiedflow], We designed three approaches that modify the self-attention operation of the velocity prediction network $\boldsymbol{v}_{\theta}$ during the denoising process. In our setup, we focus exclusively on the single transformer blocks, as they integrate information from both the source content and the conditioning input through unified modulation. While the double blocks in the underlying architecture process text and audio features separately, the single blocks concatenate these modalities, making them more suitable for controlled feature sharing. This design choice enables the model to leverage joint representations effectively, thereby enhancing its ability to preserve the structural and semantic characteristics of the source sample during generation.
@@ IV-D Classifier-Free Guidance
where $s$ denotes the CFG scale controlling the influence of the conditioning signal $y$. Higher $s$ values amplify the semantic conditioning, whereas lower values prioritize fidelity to the source. In our experiments, we employed classifier-free guidance (CFG) to control the conditioning strength during both the inversion and denoising processes. The base model, Flux that Plays Music [fluxplaysmusic], was trained with a fixed negative prompt of ”low quality, gentle”. Consequently, we adopted the same negative prompt across all our experiments to maintain consistency with the model’s training distribution. Attempts to introduce alternative negative prompts resulted in degraded performance. For instance, during timbre transfer, we used ”A recording of target instrument” and ”A recording of source instrument” as positive and negative prompts, respectively, during the denoising following the approach proposed in [adapter]. However, this configuration failed to yield meaningful results, likely due to the model’s reliance on its original negative prompt during training.
For the inversion process, we set the CFG scale to 11, whereas the model’s default value during generation is 77. Using a high CFG value (e.g., 77) in inversion was observed to push latent representations into regions of the latent space that are difficult to guide back to meaningful and attribute-aligned states, thereby hindering effective editing while preserving melody and rhythm. Conversely, during the denoising stage, we increased the CFG scale to 2020 to strongly emphasize the new conditioning prompt, ensuring that the model incorporated the desired semantic changes while maintaining musical coherence.
@@
TABLE I: The objective evaluation results on the timbre transfer.
MusicGen Supervised .220 .757 .489 .274 5.320
AudioLDM2 Zero-shot .235 .820 .527 .563 3.574
Zeta Zero-shot .224 .813 .518 .560 5.693
FluxMusic Zero-shot .220 .756 .488 .464 5.403
MusRec K Injection(ours) Zero-shot .262 .718 .490 .366 7.018
MusRec KV Injection(ours) Zero-shot .237 .851 .543 .600 4.265
MusRec V Injection(ours) Zero-shot .236 .843 .535 .583 4.605
TABLE II: The objective evaluation results on the genre transfer.
MusicGen Supervised .454 .754 .604 .129 9.790
AudioLDM2 Zero-shot .585 .698 .641 .153 8.782
Zeta Zero-shot .531 .762 .646 .315 7.158
FluxMusic Zero-shot .524 .771 .647 .354 6.774
MusRec K Injection(ours) Zero-shot .547 .754 .650 .225 11.398
MusRec KV Injection(ours) Zero-shot .545 .797 .671 .424 5.433
MusRec V Injection(ours) Zero-shot .537 .799 .668 .433 5.662
TABLE III: The subjective evaluation results on the timbre transfer.
Model MOS-T $\uparrow$ MOS-P $\uparrow$ Overall $\uparrow$
AudioLDM2 3.10 3.33 3.21
MusicGen 3.33 2.62 2.98
Zeta 3.57 3.57 3.57
FluxMusic 3.43 3.71 3.57
MusRec KV Injection (ours) 4.05 4.14 4.10
MusRec V Injection (ours) 3.90 4.24 4.07
MusRec K Injection (ours) 3.43 3.05 3.24
TABLE IV: The subjective evaluation results on the genre transfer.
Model MOS-T $\uparrow$ MOS-P $\uparrow$ Overall $\uparrow$
AudioLDM2 3.14 1.86 2.50
MusicGen 2.67 2.57 2.62
Zeta 2.71 3.76 3.24
FluxMusic 2.95 3.62 3.29
MusRec KV Injection (ours) 3.14 4.14 3.64
MusRec V Injection (ours) 3.14 4.19 3.67
MusRec K Injection (ours) 2.95 3.19 3.07
@@ A-B Full Subjective Results
Professional musicians generally assigned higher MOS-T and MOS-P ratings for models that preserved timbral detail and musical structure, showing greater sensitivity to subtle artifacts or tonal imbalances. They consistently preferred MusRec KV Injection, which provided the most faithful timbral transfer in both tasks, and rated MusRec V Injection highly for its perceptual smoothness and realism. Among the baseline systems, FluxMusic and Zeta received comparatively better scores from professionals, reflecting their stronger structural consistency, while AudioLDM2 achieved a slightly higher MOS-T in genre transfer, suggesting that its tonal balance was appreciated despite its lower perceptual realism. In contrast, MusicGen was rated the lowest due to audible artifacts and weaker genre adherence.
Ordinary listeners, on the other hand, tended to favor models that maintained overall musical coherence and recognizable style, even when minor distortions were present. For this group, MusRec V Injection often received the highest perceptual ratings, as its outputs were smoother and easier to follow, while MusRec KV Injection ranked slightly lower but remained among the top performers. Ordinary listeners also rated FluxMusic and Zeta relatively well among the baselines, likely because these models produced sonically appealing and stylistically stable results, whereas AudioLDM2 and MusicGen were perceived as less consistent.

### v2

@@ Abstract
Music editing has emerged as an important and practical area of artificial intelligence, with applications ranging from video game and film music production to personalizing existing tracks according to user preferences. However, existing models face significant limitations, such as being restricted to editing synthesized music generated by their own models, requiring highly precise prompts, or necessitating task-specific retraining—thus lacking true zero-shot capability. leveraging recent advances in rectified flow and diffusion transformers, we introduce MusRec, a zero-shot text-to-music editing model capable of performing diverse editing tasks on real-world music efficiently and effectively. Experimental results demonstrate that our approach outperforms existing methods in preserving musical content, structural consistency, and editing fidelity, establishing a strong foundation for controllable music editing in real-world scenarios.
@@ I Introduction
At the same time, a parallel line of research has introduced rectified flow models [lee2024improvingtrainingrectifiedflows, liu2022flowstraightfastlearning], which reformulate diffusion as a more direct flow between noise and data distributions. Rectified flows enable efficient and stable generation, and have recently been realized at scale through the Flux family of diffusion transformers [tang2024hart, xie2024sana, yang2024cogvideox, peebles2023scalable]. FLUX that Plays Music [fluxplaysmusic] in particular demonstrated the power of this approach for text-to-music generation. In computer vision, the work of Taming Rectified Flow for Inversion and Editing [tamingrectifiedflow] showed that RF models also support accurate inversion and robust editing, but these ideas have not yet been applied to music. This raises an intriguing question: can the strengths of rectified flow be used not just for generating music, but for editing real recordings in a practical, zero-shot fashion?
@@ I-B Contributions
  • We present rectified-flow based framework for editing of real music recordings.
@@ II-B Music Editing
In this work, we introduce MusRec, a zero-shot framework built on pretrained rectified flow models for music editing. The overall process is illustrated in  1. MusRec injects the self-attention features of the source music from a diffusion transformer directly into the editing process. Unlike prior approaches, it can operate effectively on real-world audio and generalizes across a wide range of editing tasks. Moreover, MusRec removes the need for prompt engineering during both reversal and editing, making music editing more accessible and practical.
@@ IV-B RF-Solver
$\quad+\tfrac{1}{2}(t_{i-1}-t_{i})^{2}\,\boldsymbol{v}_{\theta}^{(1)}(\boldsymbol{Z}_{t_{i}},t_{i}).$ (8)
@@ IV-C Attention Feature Replacement Strategies
To explore the effect of feature-level guidance during the denoising process, several attention feature replacement strategies within the diffusion transformer architecture were further experimented with. The goal of these experiments is to investigate how reusing intermediate representations from the inversion stage can improve controllability and structure preservation in the generated outputs. Inspired by prior work on attention-level feature reuse in rectified flow models [tamingrectifiedflow], three approaches were designed to modify the self-attention operation of the velocity prediction network $\boldsymbol{v}_{\theta}$ during the denoising process. In this setup, the focus is exclusively on the single transformer blocks, as they integrate information from both the source content and the conditioning input through unified modulation. While the double blocks in the underlying architecture process text and audio features separately, the single blocks concatenate these modalities, making them more suitable for controlled feature sharing. This design choice enables the model to leverage joint representations effectively, thereby enhancing its ability to preserve the structural and semantic characteristics of the source sample during generation.
@@ IV-D Classifier-Free Guidance
where $s$ denotes the CFG scale controlling the influence of the conditioning signal $y$. Higher $s$ values amplify the semantic conditioning, whereas lower values prioritize fidelity to the source. In the experiments, classifier-free guidance (CFG) was employed to control the conditioning strength during both the inversion and denoising processes. The base model, FLUX that Plays Music [fluxplaysmusic], was trained with a fixed negative prompt of ”low quality, gentle”. Consequently, the same negative prompt was adopted across all experiments to maintain consistency with the model’s training distribution. Attempts to introduce alternative negative prompts resulted in degraded performance. For instance, during timbre transfer, ”A recording of target instrument” and ”A recording of source instrument” were used as positive and negative prompts during the denoising [adapter]. However, this configuration failed to yield meaningful results, likely due to the model’s reliance on its original negative prompt during training.
For the inversion process, the CFG scale was set to $1$, whereas the model’s default value during generation is $7$. Using a high CFG value (e.g., $7$) in inversion was observed to push latent representations into regions of the latent space that are difficult to guide back to meaningful and attribute-aligned states, thereby hindering effective editing while preserving melody and rhythm. Conversely, during the denoising stage, the CFG scale was increased to $20$ to strongly emphasize the new conditioning prompt, ensuring that the model incorporated the desired semantic changes while maintaining musical coherence.
@@ V-D Subjective Metrics
TABLE I: The objective evaluation results on the timbre transfer.
MusicGen Supervised .220 .757 .489 .274 5.320
AudioLDM2 Zero-shot .235 .820 .527 .563 3.574
Zeta Zero-shot .224 .813 .518 .560 5.693
FluxMusic Zero-shot .220 .756 .488 .464 5.403
MusRec K Injection(ours) Zero-shot .262 .718 .490 .366 7.018
MusRec KV Injection(ours) Zero-shot .237 .851 .543 .600 4.265
MusRec V Injection(ours) Zero-shot .236 .843 .535 .583 4.605
TABLE II: The objective evaluation results on the genre transfer.
MusicGen Supervised .454 .754 .604 .129 9.790
AudioLDM2 Zero-shot .585 .698 .641 .153 8.782
Zeta Zero-shot .531 .762 .646 .315 7.158
FluxMusic Zero-shot .524 .771 .647 .354 6.774
MusRec K Injection(ours) Zero-shot .547 .754 .650 .225 11.398
MusRec KV Injection(ours) Zero-shot .545 .797 .671 .424 5.433
MusRec V Injection(ours) Zero-shot .537 .799 .668 .433 5.662
@@ VI-B Objective Results
TABLE III: The subjective evaluation results on the timbre transfer.
Model MOS-T $\uparrow$ MOS-P $\uparrow$ Overall $\uparrow$
AudioLDM2 3.10 3.33 3.21
MusicGen 3.33 2.62 2.98
Zeta 3.57 3.57 3.57
FluxMusic 3.43 3.71 3.57
MusRec KV Injection (ours) 4.05 4.14 4.10
MusRec V Injection (ours) 3.90 4.24 4.07
MusRec K Injection (ours) 3.43 3.05 3.24
TABLE IV: The subjective evaluation results on the genre transfer.
Model MOS-T $\uparrow$ MOS-P $\uparrow$ Overall $\uparrow$
AudioLDM2 3.14 1.86 2.50
MusicGen 2.67 2.57 2.62
Zeta 2.71 3.76 3.24
FluxMusic 2.95 3.62 3.29
MusRec KV Injection (ours) 3.14 4.14 3.64
MusRec V Injection (ours) 3.14 4.19 3.67
MusRec K Injection (ours) 2.95 3.19 3.07
@@ A-B Full Subjective Results
Professional musicians generally assigned higher MOS-T and MOS-P ratings for models that preserved timbral detail and musical structure, showing greater sensitivity to subtle artifacts or tonal imbalances. They consistently preferred MusRec KV Injection, which provided the most faithful timbral transfer in both tasks, while MusRec V Injection also received high ratings from participants. Among the baseline systems, FluxMusic and Zeta received comparatively better scores from professionals, while AudioLDM2 achieved a slightly higher MOS-T in genre transfer, suggesting that its tonal balance was appreciated despite its lower perceptual realism. In contrast, MusicGen was rated the lowest among all models.
Ordinary listeners, on the other hand, tended to favor models that maintained overall musical coherence and recognizable style, even when minor distortions were present. For this group, MusRec V Injection often received the highest perceptual ratings, as its outputs were smoother and easier to follow, while MusRec KV Injection ranked slightly lower but remained among the top performers. Ordinary listeners also rated FluxMusic and Zeta relatively well among the baselines, whereas AudioLDM2 and MusicGen were perceived as less consistent.